# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v3

@@ 3 A Proxy to the Joint Sampling Distribution
To motivate our design methods proposed in Section 4, we develop a proxy to the joint sampling distribution of posterior probabilities. These proxies are required for the theory that substantiates our methodology, but our methods do not directly use them. Instead, we estimate the true joint sampling distribution of posterior probabilities by generating data $\boldsymbol{y}^{{}_{(n)}}$ using the straightforward process in Algorithm 1. Our proxies leverage the regularity conditions listed in Appendix A of the supplement. Appendix A.1 details the four necessary assumptions to invoke the Bernstein-von Mises (BvM) theorem (van der Vaart, 1998). The first three assumptions are weaker than the regularity conditions for the asymptotic normality of the maximum likelihood estimator (MLE) (Lehmann and Casella, 1998), which are listed in Appendix A.2.
@@
where $\hat{\boldsymbol{\theta}}^{{}_{(n)}}_{r}$ is the maximum likelihood estimate of $\boldsymbol{\theta}$, $\mathcal{I}(\cdot)$ is the Fisher information with respect to the targets of inference, and $\boldsymbol{\theta}_{r}=\{\theta_{r,k}\}_{k=1}^{K}=\{g_{k}(\boldsymbol{\eta}_{r})\}_{k=1}^{K}$. We approximate the joint posterior of $\boldsymbol{\theta}$ in (2) to allow for an arbitrary dependence structure between its components. Even if all components of $\boldsymbol{\theta}$ are modeled independently, a joint posterior can still be constructed. The approximate sampling distribution of the MLE $\hat{\boldsymbol{\theta}}^{{}_{(n)}}~{}|~{}\boldsymbol{\eta}=\boldsymbol{\eta}_{r}$ is $\mathcal{N}(\boldsymbol{\theta}_{r},n^{-1}\mathcal{I}(\boldsymbol{\theta}_{r})^{-1})$ under the regularity conditions in Lehmann and Casella (1998). To develop our proxy used for theoretical purposes, a single realization from this $K$-dimensional multivariate normal distribution could be generated using conditional cumulative distribution function (CDF) inversion and a point $\boldsymbol{u}=\{u_{k}\}_{k=1}^{K}\in[0,1]^{K}$. We could obtain the first component $\hat{\theta}^{{}_{(n)}}_{r,1}$ as the $u_{1}$-quantile of the sampling distribution of $\hat{\theta}^{{}_{(n)}}_{1}~{}|~{}\boldsymbol{\eta}_{r}$. For the remaining components, we could generate $\hat{\theta}^{{}_{(n)}}_{r,k}$ as the $u_{k}$-quantile of the sampling distribution of $\hat{\theta}^{{}_{(n)}}_{k}~{}|~{}\{\hat{\theta}^{{}_{(n)}}_{s}=\hat{\theta}^{{}_{(n)}}_{r,s}\}_{s=1}^{k-1},\hskip 0.1pt\boldsymbol{\eta}_{r}$.
@@
where $\Phi(\cdot)$ is the standard normal CDF. The collection of $\{p^{{}_{(n)}}_{r,k}\}_{k=1}^{K}$ values corresponding to $\{\boldsymbol{u}_{r}\}_{r=1}^{m}\sim\mathcal{U}\left([0,1]^{K}\right)$ and $\{\boldsymbol{\eta}_{r}\}_{r=1}^{m}\sim\Psi$ define our proxy to the joint sampling distribution of posterior probabilities. Because this proxy to the sampling distribution relies on asymptotic results, it may differ materially from the true sampling distribution of posterior probabilities for finite $n$.
Thus, the proxy sampling distribution only motivates our theoretical result in Theorem 1. This result guarantees that the logit of $p^{{}_{(n)}}_{r,k}$ is an approximately linear function of $n$ for each metric $k$. We later adapt this result to assess the operating characteristics of A/B tests across a wide range of sample sizes by estimating the true joint sampling distribution of posterior probabilities under $\Psi$ at only two values of $n$. Each $p^{{}_{(n)}}_{r,k}$ value depends on the value for $\hat{\boldsymbol{\theta}}^{{}_{(n)}}_{r}$, which in turn depends on the sample size $n$, the parameter value $\boldsymbol{\eta}_{r}$, and the point $\boldsymbol{u}_{r}$. In Theorem 1, we fix both $\boldsymbol{\eta}_{r}$ and $\boldsymbol{u}_{r}$ to examine the behaviour of $p^{{}_{(n)}}_{r,k}$ as a deterministic function of $n$.
@@
For any $\boldsymbol{\eta}_{r}\sim\Psi$, let the model $f(\boldsymbol{y};\boldsymbol{\eta}_{r})$ satisfy the conditions in Appendix A.2 and the prior $p(\boldsymbol{\eta})$ satisfy the conditions in Appendix A.1. Define $\emph{logit}(x)=\emph{log}(x)-\emph{log}(1-x)$. We consider a given point $\boldsymbol{u}_{r}\in[0,1]^{K}$. For $k=1,\dots,K$, the functions $p^{{}_{(n)}}_{r,k}$ in (3) are such that
@@
    $p^{{}_{(n)}}_{r,k}=\Phi\left(a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})\sqrt{n}+b_{k}(\boldsymbol{u}_{r})\right)-\Phi\left(a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})\sqrt{n}+b_{k}(\boldsymbol{u}_{r})\right)$, where $a_{k}(\cdot)$ and $b_{k}(\cdot)$ are functions that do not depend on $n$.
@@
    $\lim\limits_{n\rightarrow\infty}\dfrac{d}{dn}~{}\emph{logit}\left(p^{{}_{(n)}}_{r,k}\right)=(0.5-\mathbb{I}\{\theta_{r,k}\notin(\delta_{L,k},\delta_{U,k})\})\times\emph{min}\{a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})^{2},a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})^{2}\}$.
We prove parts $(a)$ and $(b)$ of Theorem 1 in Appendix B of the supplement. Hagar and Stevens (2025) developed similar though simplified theory for a single hypothesis, and Hagar and Golchi (2025) derived a result similar to Theorem 1 that indexed the proxy sampling distribution by a function of the sample size after adjusting the posterior to mitigate confounding in clinical settings. However, neither of those previous results can be directly used to design A/B tests while controlling the FDR and average power. We now consider the practical implications of Theorem 1. The limiting derivative in part $(b)$ is a constant that does not depend on $n$. For the probabilities corresponding to each metric in the joint sampling distribution, the linear approximation to $l^{{}_{(n)}}_{r,k}=\text{logit}(p^{{}_{(n)}}_{r,k})$ as a function of $n$ is a good global approximation for sufficiently large sample sizes. Moreover, this linear approximation should be locally suitable for a range of sample sizes.
It follows that the quantiles of the sampling distribution of $l^{{}_{(n)}}_{r,k}$ change linearly as a function of $n$ when $\boldsymbol{\theta}_{r}$ is held constant across simulation repetitions. In Section 4, we exploit and adapt this linear trend in the proxy sampling distribution to flexibly model the logits of posterior probabilities as linear functions of $n$ when independently simulating samples $\boldsymbol{y}^{{}_{(n)}}$ according to $\boldsymbol{\theta}_{r}\sim\Psi$ as in Algorithm 1. To ensure our method performs well with finite sample sizes, we only use the limiting slopes from Theorem 1 to initialize our approach. The good performance of our simulation-based method, which empirically estimates linear functions from the data, is illustrated in Section 5.
@@ 4 Economical Assessment of the FDR and Average Power
We generalize the results from Theorem 1 to develop the procedure in Algorithm 2. This procedure allows practitioners to efficiently explore the ($n,\gamma$)-space to find the $(n,\gamma)$ combination that minimizes the sample size while satisfying criteria for the FDR and average power. Algorithm 2 is economical because we estimate the true joint sampling distribution of posterior probabilities via simulation at only two sample sizes: $n_{0}$ and $n_{1}$. The initial sample size $n_{0}$ is an input for Algorithm 2, and it can be selected based on the anticipated budget or timeline for the A/B test. We note that Algorithm 2 details a general application of our methodology, and we later describe potential modifications.
Algorithm 2 Procedure to Determine Optimal Sample Size and Decision Threshold
@@
2:     Compute $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{0})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ as in Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{0})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
@@
6:     Find $n_{1}$, the smallest $n$ such that Algorithm 1 with the inverse logits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average      power $\geq 1-\beta$, where $\gamma_{n}$ is the smallest probability such that $\widehat{\text{FDR}}$ $\leq q$
7:     Compute $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{1})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ as in Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
@@
13:     Find $n_{2}$, the smallest $n$ such that Algorithm 1 with the inverse logits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average      power $\geq 1-\beta$, where $\gamma_{n}$ is the smallest probability such that $\widehat{\text{FDR}}$ $\leq q$
14:     return $n_{2}$ as recommended $n$ and $\gamma_{n_{2}}$ as $\gamma$
@@
We now elaborate on several steps of Algorithm 2. In Line 2, we estimate posterior probabilities at the sample size $n_{0}$ by simulating data as in Algorithm 1. These posterior probabilities can flexibly be computed using any computational or analytical approximation method. If a computational method is used to generate posterior samples, we recommend calculating posterior probabilities using a nonparametric kernel density estimate of the posterior so that the logits of all probabilities are finite. The notation $\hat{l}^{{}_{(n)}}_{r,k}$ is also introduced in Line 2. These estimated logits from the true joint sampling distribution leverage independently generated samples $\boldsymbol{y}^{{}_{(n)}}_{r}$ from $\Psi$ for each simulation repetition $r$. Unlike for $l^{{}_{(n)}}_{r,k}$ from the proxy sampling distribution in Theorem 1, there is no relationship between the $\hat{l}^{{}_{(n)}}_{r,k}$ values corresponding to two different sample sizes that happen to have the same indices for $r$ and $k$. In Line 5 of Algorithm 2, we construct linear approximations to the logits of posterior probabilities as a function of $n$ using the limiting slopes from Theorem 1. For moderate $n$, the limiting slopes for $l^{{}_{(n)}}_{r,k}$ may not be accurate since Theorem 1 relies on the approximate normality of the posterior that is asymptotically guaranteed when the conditions for the BvM theorem are satisfied. Thus, we only use those slopes to initialize our method.
@@
We choose the next sample size $n_{1}$ at which to estimate posterior probabilities in Line 6 of Algorithm 2. Recall that the optimal decision threshold $\gamma$ generally decreases with $n$ as described in Section 2. To account for this phenomenon, we let $\gamma$ be the smallest probability for which $\widehat{\text{FDR}}$, the estimate of the FDR defined in (1), is at most $q$. Line 6 introduces a subscript $n$ for the decision threshold to emphasize this dependence on the sample size. Given $\gamma_{n}$ for a particular $n$, we can determine whether this sample size achieves an average power of at least $1-\beta$. The sample size $n_{1}$ is the smallest $n$ such that the average power is sufficiently large. We recommend finding this sample size using binary search. We emphasize that a modified version of Algorithm 1 based on the linear approximations from Line 5 of Algorithm 2 can be quickly implemented for new sample sizes without computing additional posterior probabilities.
@@
In Lines 10 to 12 of Algorithm 2, we construct linear approximations to logits of posterior probabilities that are less reliant on large-sample results. These approximations are obtained separately for each metric $k$ by fitting a line through the same order statistic of the marginal sampling distribution at the sample sizes $n_{0}$ and $n_{1}$. To maintain the proper level of dependence in the joint sampling distribution of posterior probabilities, we group the lines $\hat{L}^{{}_{(n)}}_{r,k}$ from Line 12 across metrics based on the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ that defines the linear approximations. Given the linear trend in the proxy sampling distribution quantiles discussed in Section 3, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the $\boldsymbol{\theta}_{r}$ value is similar for all $\boldsymbol{\eta}_{r}\sim\Psi$. However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 12 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision threshold in Line 13.
Lastly, we quantify the impact of simulation variability by constructing bootstrap confidence intervals for the optimal sample size and decision threshold. We construct these confidence intervals by sampling $m$ times with replacement from $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{0})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ and $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{1})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ obtained in Algorithm 2. We note that posterior probabilities for each of the $K$ individual metrics are resampled jointly. If $\Psi$ is comprised of various submodels, this resampling process should be implemented separately for each submodel to facilitate constructing new linear approximations. We obtain a new $(n,\gamma)$ recommendation by implementing the process in Lines 8 to 13 of Algorithm 2 with the bootstrap estimates of the sampling distributions at $n_{0}$ and $n_{1}$. This process is repeated $M$ times, and bootstrap confidence intervals for the optimal $n$ and $\gamma$ values are calculated using the percentile method (Efron, 1982). The width of these confidence intervals can help inform the choice for the number of simulation repetitions $m$. The suitable coverage properties of such confidence intervals were confirmed in Hagar and Stevens (2025). In Section 5, we consider the performance of Algorithm 2 and construct bootstrap confidence intervals for an example based on a real A/B test involving 5 metrics.

### v4

@@ 3 A Proxy to the Joint Sampling Distribution
To motivate our design methods proposed in Section 4, we develop a proxy to the joint sampling distribution of posterior probabilities. These proxies are required for the theory that substantiates our methodology, but our methods do not directly use them. Instead, we estimate the true joint sampling distribution of posterior probabilities by generating data $\boldsymbol{y}^{{}_{(n)}}$ using the straightforward process in Algorithm 1. Our proxies leverage the regularity conditions listed in Appendix A of the supplement. Appendix A.1 details the four necessary assumptions to invoke the Bernstein-von Mises (BvM) theorem (van der Vaart, 1998). The first three assumptions are also required for the asymptotic normality of the maximum likelihood estimator (MLE) (van der Vaart, 1998), and the MLE regularity conditions are listed in Appendix A.2.
@@
where $\hat{\boldsymbol{\theta}}^{{}_{(n)}}_{r}$ is the maximum likelihood estimate of $\boldsymbol{\theta}$, $\mathcal{I}(\cdot)$ is the Fisher information with respect to the targets of inference, and $\boldsymbol{\theta}_{r}=\{\theta_{r,k}\}_{k=1}^{K}=\{g_{k}(\boldsymbol{\eta}_{r})\}_{k=1}^{K}$. We approximate the joint posterior of $\boldsymbol{\theta}$ in (4) to allow for an arbitrary dependence structure between its components. The approximate sampling distribution of the MLE $\hat{\boldsymbol{\theta}}^{{}_{(n)}}~{}|~{}\boldsymbol{\eta}=\boldsymbol{\eta}_{r}$ is $\mathcal{N}(\boldsymbol{\theta}_{r},n^{-1}\mathcal{I}(\boldsymbol{\theta}_{r})^{-1})$ under the regularity conditions in van der Vaart (1998). To develop our proxy used for theoretical purposes, a single realization from this $K$-dimensional multivariate normal distribution could be generated using conditional cumulative distribution function (CDF) inversion and a point $\boldsymbol{u}=\{u_{k}\}_{k=1}^{K}\in[0,1]^{K}$. We could obtain the first component $\hat{\theta}^{{}_{(n)}}_{r,1}$ as the $u_{1}$-quantile of the sampling distribution of $\hat{\theta}^{{}_{(n)}}_{1}~{}|~{}\boldsymbol{\eta}_{r}$. For the remaining components, we could iteratively generate $\hat{\theta}^{{}_{(n)}}_{r,k}$ as the $u_{k}$-quantile of the sampling distribution of $\hat{\theta}^{{}_{(n)}}_{k}~{}|~{}\{\hat{\theta}^{{}_{(n)}}_{s}=\hat{\theta}^{{}_{(n)}}_{r,s}\}_{s=1}^{k-1},\hskip 0.1pt\boldsymbol{\eta}_{r}$.
@@
where $\Phi(\cdot)$ is the standard normal CDF. The collection of $\{p^{{}_{(n)}}_{r,k}\}_{k=1}^{K}$ values corresponding to $\{\boldsymbol{u}_{r}\}_{r=1}^{m}\sim\mathcal{U}\left([0,1]^{K}\right)$ and $\{\boldsymbol{\eta}_{r}\}_{r=1}^{m}\sim\Psi$ define our proxy to the joint sampling distribution of posterior probabilities. Theory from Hagar and Stevens (2024) can be extended to show that the total variation distance between the proxy sampling distribution and the true sampling distribution converges in probability to 0 as $n\rightarrow\infty$. However, the proxy and true sampling distributions could differ materially for finite $n$.
Thus, the proxy sampling distribution only motivates our theoretical result in Theorem 1. This result guarantees that the logit of $p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}$ is an approximately linear function of $n$ for each metric $k$. For $k=1$, the conditioning set $\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}$ is empty. We later adapt this result to assess the operating characteristics of A/B tests across a wide range of sample sizes by estimating the true joint sampling distribution of posterior probabilities under $\Psi$ at only two values of $n$. Each $p^{{}_{(n)}}_{r,k}$ value depends on the value for $\hat{\boldsymbol{\theta}}^{{}_{(n)}}_{r}$, which in turn depends on the sample size $n$, the parameter value $\boldsymbol{\eta}_{r}$, and the point $\boldsymbol{u}_{r}$. In Theorem 1, we fix both $\boldsymbol{\eta}_{r}$ and $\boldsymbol{u}_{r}$ to examine the behaviour of $p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}$ as a deterministic function of $n$.
@@
For any $\boldsymbol{\eta}_{r}\sim\Psi$, let the model $f(\boldsymbol{y};\boldsymbol{\eta}_{r})$ satisfy the conditions in Appendix A.2 and the prior $p(\boldsymbol{\eta})$ satisfy the conditions in Appendix A.1. Define $\emph{logit}(x)=\emph{log}(x)-\emph{log}(1-x)$. We consider a given point $\boldsymbol{u}_{r}=\{u_{r,k}\}_{k=1}^{K}\in[0,1]^{K}$. For $k=1,\dots,K$, the functions $p^{{}_{(n)}}_{r,k}$ in (5) are such that
@@
    $p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}=\Phi\left(a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})\sqrt{n}+b_{k}(\{u_{r,h}\}_{h=1}^{k})\right)-\Phi\left(a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})\sqrt{n}+b_{k}(\{u_{r,h}\}_{h=1}^{k})\right)$, where $a_{k}(\delta_{k},\boldsymbol{\theta}_{r})=(\delta_{k}-\theta_{r,k})/\sqrt{\mathcal{I}_{k,k}(\boldsymbol{\theta}_{r})^{-1}}$ and $b_{k}(\cdot)$ is a not a function of $n$.
@@
    $\lim\limits_{n\rightarrow\infty}\dfrac{d}{dn}~{}\emph{logit}\left(p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}\right)=(0.5-\mathbb{I}\{\theta_{r,k}\notin(\delta_{L,k},\delta_{U,k})\})\times\emph{min}\{a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})^{2},a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})^{2}\}$.
We prove parts $(a)$ and $(b)$ of Theorem 1 in Appendix B of the supplement. Hagar and Stevens (2025) developed simplified theory for $p^{{}_{(n)}}_{r,k}$ corresponding to a single hypothesis. However, that result does not account for the dependence structure across multiple hypotheses and cannot be used to model joint sampling distributions of posterior probabilities, hence the need for the extended theory. We now consider the practical implications of Theorem 1. The limiting derivative in part $(b)$ is a constant that does not depend on $n$. Moreover, these limiting derivatives do not depend on the point $\boldsymbol{u}_{r}$, which controls the dependence in the joint proxy sampling distribution. The limiting derivatives for $\text{logit}\left(p^{{}_{(n)}}_{r,k}\right)$ and $\text{logit}\left(p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}\right)$ are therefore the same. For the probabilities corresponding to each metric in the joint sampling distribution, the linear approximation to $l^{{}_{(n)}}_{r,k}=\text{logit}(p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1})$ as a function of $n$ is a good global approximation for sufficiently large sample sizes. Moreover, this linear approximation should be locally suitable for a range of sample sizes.
It follows that the (conditional) quantiles of the sampling distribution of $l^{{}_{(n)}}_{r,k}$ change linearly as a function of $n$ when $\boldsymbol{\theta}_{r}$ is held constant across simulation repetitions. In Section 4, we exploit and adapt this linear trend in the proxy sampling distribution to flexibly model the logits of posterior probabilities as linear functions of $n$ when independently simulating samples $\boldsymbol{y}^{{}_{(n)}}$ according to $\boldsymbol{\theta}_{r}\sim\Psi$ as in Algorithm 1. To ensure our method performs well with finite sample sizes, we only use the limiting slopes from Theorem 1 to initialize our approach. The good performance of our simulation-based method, which empirically estimates linear functions from the data, is illustrated in Section 5.
@@ 4 Economical Assessment of the FDR and Average Power
We generalize the results from Theorem 1 to develop the procedure in Algorithm 2. This procedure allows practitioners to efficiently explore the ($n,\boldsymbol{\gamma}$)-space to find the $(n,\boldsymbol{\gamma})$ combination under a given optimization scheme that minimizes the sample size while satisfying criteria for the FDR and average power. Algorithm 2 is economical because we estimate the true joint sampling distribution of posterior probabilities via simulation at only two sample sizes: $n_{0}$ and $n_{1}$. The initial sample size $n_{0}$ is an input for Algorithm 2, and it can be selected based on the anticipated budget or timeline for the A/B test. We note that Algorithm 2 details a general application of our methodology, and we later describe potential modifications.
Algorithm 2 Procedure to Determine Optimal Sample Size and Decision Thresholds
@@
2:     Compute $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{0})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ via Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{0})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
@@
6:     Find $n_{1}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power      $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
7:     Compute $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{1})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ via Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
@@
13:     Find $n_{2}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power      $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
14:     return $n_{2}$ as recommended $n$ and $\boldsymbol{\gamma}_{n_{2}}$ as $\boldsymbol{\gamma}$
@@
We now elaborate on several steps of Algorithm 2. In Line 2, we estimate posterior probabilities at the sample size $n_{0}$ by simulating data as in Algorithm 1. These posterior probabilities can flexibly be computed using any computational or analytical approximation method. If a computational method is used to generate posterior samples, we recommend calculating posterior probabilities using a nonparametric kernel density estimate of the posterior so that the logits of all probabilities are finite. The notation $\hat{l}^{{}_{(n)}}_{r,k}$ is also introduced in Line 2. These estimated logits from the true joint sampling distribution leverage independently generated samples $\boldsymbol{y}^{{}_{(n)}}_{r}$ from $\Psi$ for each simulation repetition $r$. Because we estimate the joint sampling distribution of posterior probabilities, the resulting estimated logits allow us to model logits of the conditional posterior probabilities considered in Section 3.
Unlike for $l^{{}_{(n)}}_{r,k}$ from the proxy sampling distribution in Theorem 1, there is no relationship between the $\hat{l}^{{}_{(n)}}_{r,k}$ values corresponding to two different sample sizes that happen to have the same indices for $r$ and $k$. In Line 5 of Algorithm 2, we construct linear approximations to the logits of posterior probabilities as a function of $n$ using the limiting slopes from Theorem 1. For moderate $n$, the limiting slopes for $l^{{}_{(n)}}_{r,k}$ may not be accurate since Theorem 1 relies on asymptotic theory. Thus, we only use those slopes to initialize our method.
@@
We choose the next sample size $n_{1}$ at which to estimate posterior probabilities in Line 6 of Algorithm 2. Recall that the optimal decision thresholds $\boldsymbol{\gamma}$ are linked with $n$ as described in Section 2. To account for this phenomenon, we run an optimization scheme to find the optimal $\boldsymbol{\gamma}$ for which $\widehat{\text{FDR}}$, the estimate of the FDR defined in (1), is at most $q$. This optimization scheme can be as simple or as complex as the practitioner desires; we elaborate on several potential schemes in Section 5. Line 6 introduces a subscript $n$ for the decision thresholds to emphasize this dependence on the sample size. Given $\boldsymbol{\gamma}_{n}$ for a particular $n$, we can determine whether this sample size achieves an average power of at least $1-\beta$. The sample size $n_{1}$ is the smallest $n$ such that the average power is sufficiently large. We recommend finding this sample size using binary search. We emphasize that a modified version of Algorithm 1 based on the linear approximations from Line 5 of Algorithm 2 can be efficiently implemented for new sample sizes without computing additional posterior probabilities.
@@
In Lines 10 to 12 of Algorithm 2, we construct linear approximations to logits of posterior probabilities that are less reliant on large-sample results. These approximations are obtained separately for each metric $k$ by fitting a line through the same order statistic of the marginal sampling distribution at the sample sizes $n_{0}$ and $n_{1}$. Because Theorem 1 ensures that the limiting slopes of $\text{logit}\left(p^{{}_{(n)}}_{r,k}\right)$ and $\text{logit}\left(p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}\right)$ from the proxy sampling distribution are the same, we use the marginal sampling distributions to estimate slopes for the conditional logits. To maintain the proper level of dependence in the joint sampling distribution of posterior probabilities, we group the lines $\hat{L}^{{}_{(n)}}_{r,k}$ from Line 12 across metrics based on the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ that defines the linear approximations. Given the linear trend in the proxy sampling distribution quantiles discussed in Section 3, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the $\boldsymbol{\theta}_{r}$ value is similar for all $\boldsymbol{\eta}_{r}\sim\Psi$.
However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 12 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision thresholds in Line 13. As a result of Theorem 1 and the total variation distance between the proxy and true sampling distributions converging to 0 as $n\rightarrow\infty$, Algorithm 2 guarantees the FDR control in (2). In Section 5, we consider the performance of this algorithm for an example based on a real A/B test involving 5 metrics.